# Differentially Private Adaptation of Diffusion Models via Noisy Aggregated Embeddings

### v1

@@ Abstract
We introduce novel methods for adapting diffusion models under differential privacy (DP) constraints, enabling privacy-preserving style and content transfer without fine-tuning. Traditional approaches to private adaptation, such as DP-SGD, incur significant computational overhead and degrade model performance when applied to large, complex models. Our approach instead leverages embedding-based techniques: Universal Guidance and Textual Inversion (TI), adapted with differentially private mechanisms. We apply these methods to Stable Diffusion for style adaptation using two private datasets: a collection of artworks by a single artist and pictograms from the Paris 2024 Olympics. Experimental results show that the TI-based adaptation achieves superior fidelity in style transfer, even under strong privacy guarantees, while both methods maintain high privacy resilience by employing calibrated noise and subsampling strategies. Our findings demonstrate a feasible and efficient pathway for privacy-preserving diffusion model adaptation, balancing data protection with the fidelity of generated images, and offer insights into embedding-driven methods for DP in generative AI applications.
@@ 1 Introduction
In recent years, diffusion models ho2020denoising ; song2020score , particularly latent diffusion models rombach2022high , have spearheaded high quality text-to-image generation, and have been widely adopted by researchers and the general public alike. Trained on massive datasets like LAION-5B schuhmann2022laion , these models have developed a broad understanding of visual concepts, enabling new creative and practical applications. Notably, tools like Stable Diffusion rombach2022high ; podell2023sdxl have been made readily accessible for general use. Building on this foundation, efficient adaptation methods such as parameter efficient fine-tuning (PEFT) hu2022lora ; von-platen-etal-2022-diffusers ; ruiz2023dreambooth , guidance based approaches ho2021classifierfree ; kim2022diffusionclip ; bansal2023universal , and pseudo-word generation gal2023an enable users to leverage this extensive pretraining for customizing models that can specialize on downstream tasks with smaller datasets.
@@
However, the rapid adoption of diffusion models has also raised significant privacy, ethical and legal concerns. One critical issue is the vulnerability of these models to privacy attacks, from membership inference duan2023diffusion , where an attacker determines whether a specific data point was used to train a particular model, to data extraction carlini2023extracting , which enables an attacker to reconstruct particular images from the training dataset. This issue is even more severe during the fine-tuning phase where the model is fine-tuned on smaller specialized datasets from a possibly different domain and each data record has more impact on the final model. This risk underscores the importance of privacy-preserving technologies, particularly as diffusion models often rely on vast datasets scraped from the internet without explicit consent from content owners. Copyright concerns further complicate the landscape vyas2023provable . In particular, diffusion models can closely reproduce original artworks without credit or compensation to the artist, as evident in data extraction attacks carlini2023extracting . Such concerns intensify the debate on balancing technological advancement with the rights and privacy of content owners, underscoring the need for clearer ethical and legal guidelines in the training and deployment of generative models.
@@
Differential privacy (DP) dwork2006calibrating ; dwork2006differential is a widely adopted approach to addressing these challenges, where controlled noise is added during training or inference to prevent information leakage from individual data points while still enabling the model to learn effectively from the overall dataset. The standard method for ensuring DP in deep learning is Differentially Private Stochastic Gradient Descent (DP-SGD) abadi2016deep , which modifies traditional SGD by adding noise to clipped gradients.
@@
While, in theory, DP-SGD can be applied to train diffusion models, it poses several significant practical challenges. First, it is well known that DP-SGD introduces considerable computational and memory overhead due to the need for per-sample gradient clipping, which is crucial for bounding sensitivity dwork2006calibrating ; abadi2016deep . Second, DP-SGD is fundamentally incompatible with batch-wise operations, such as batch normalization, since these operations inherently link samples together, rendering sensitivity analysis infeasible. Moreover, training a large network with DP-SGD typically leads to significant performance degradation (i.e. the curse of dimensionality), even under relatively lenient privacy budgets. To the best of our knowledge, existing diffusion models trained with DP-SGD are constrained to relatively small-scale datasets dockhorn2023differentially ; ghalebikesabi2023differentially .
@@
As a result, recent research has increasingly turned to alternative strategies for adapting diffusion models in a privacy-preserving manner without the need for training from scratch under DP-SGD. A promising approach leverages large, publicly pre-trained models, adapting them to new domains under differential privacy (DP) constraints, which allows models to capitalize on existing representational strengths while avoiding the computational and memory burdens of full DP training ghalebikesabi2023differentially . Similarly, parameter-efficient fine-tuning (PEFT) methods, such as DP-LoRA yu2021differentially , fine-tune only a subset of parameters, allowing for efficient, targeted adaptation with limited privacy costs. Methods like DP-RDM (Differentially Private Retrieval-Augmented Diffusion Model) lebensold2024dp illustrate this trend by enabling model adaptation through a retrieval mechanism that conditions image generation on private data retrieved at inference time, avoiding direct model updates. These approaches collectively demonstrate how large models can be adapted privately, circumventing the complexities and high costs associated with full DP training.
Meanwhile, alternative methods such as Universal Guidance bansal2023universal and Textual Inversion (TI) gal2023an enable model adaptation to specific content or style without modifying the diffusion model itself, instead relying on external embedding vectors. Universal Guidance’s style guidance (SG) mechanism generates a target CLIP embedding by passing the reference image through a CLIP image encoder. During the reverse diffusion process, the cosine distance between this target embedding and the CLIP embedding of the generated image is minimized, guiding the generation process to match the desired style. TI, in contrast, learns a new token embedding representing a set of target images, which is then incorporated into the text prompt to adapt the model to the style or content associated with those images. Both SG and TI avoid direct model optimization, reducing memory and computation demands while also involving fewer trainable parameters, making them highly compatible with DP constraints on smaller datasets.
In this work, we propose two novel privacy-preserving adaptation methods for smaller datasets, leveraging SG and TI to bypass the need for DP-SGD or DP-PEFT approaches that require extensive model updating. Our first approach adapts Universal Guidance for privacy by using a differentially private CLIP embedding. Although SG is typically applied to a single target image, we extend it to handle multiple images under DP by computing an embedding for each target image and aggregating these embeddings via a centroid calculation. To ensure differential privacy, we add calibrated noise to the centroid, which protects individual data points in the dataset while preserving the model’s guidance capabilities.
Our second approach enhances Textual Inversion (TI) with differential privacy. While standard TI compresses training data into a single, low-dimensional vector, offering some intrinsic privacy benefits, it lacks the rigorous guarantees required by DP. We address this by introducing a private variant of TI, in which we decouple interactions among samples by learning a separate embedding for each target image. These embeddings are then aggregated into a noisy centroid, similar to our SG approach, ensuring differential privacy while preserving TI’s data compression advantages.
Our experimental results underscore the practicality and effectiveness of the proposed differentially private embedding-based adaptation methods, with Textual Inversion (TI) proving particularly robust for preserving stylistic fidelity even under privacy constraints. In applying our approach to two distinct datasets–a private collection of artworks by @eveismyname and pictograms from the Paris 2024 Olympics Paris2024Pictograms –we demonstrate that differentially private TI can successfully capture nuanced stylistic elements while maintaining privacy guarantees. Our findings reveal a trade-off between privacy levels (determined by the DP parameter $\varepsilon$) and image quality, where lower $\varepsilon$ values yield diminished stylistic fidelity, yet still preserve the target style under moderate noise. Furthermore, subsampling offers an effective means of improving privacy resilience by controlling sensitivity to individual data points, significantly mitigating the impact of added noise on image quality. Ultimately, our approach provides a versatile framework for privacy-preserving model adaptation, enabling diffusion models to generalize effectively to novel styles and domains without risking exposure of sensitive data points.
2 Background and Related Works
@@ 2.1 Diffusion Models
Diffusion models ho2020denoising ; song2020score ; song2020denoising ; rombach2022high leverage an iterative denoising process to generate high-quality images that aligns with a given conditional input from random noise. In text-to-image generation, this conditional input is based on a textual description (a prompt) that guides the model in shaping the image to reflect the content and style specified by the text. To convert the text prompt into a suitable conditional format, it is first tokenized into discrete tokens, each representing a word or sub-word unit. These tokens are then converted into a sequence of embedding vectors $v_{i}$ that encapsulate the meaning of each token within the model’s semantic space. Next, these embeddings pass through a transformer text encoder, such as CLIP radford2021learning , outputting a single text-conditional vector $y$ that serves as the conditioning input. This text-conditioning vector $y$ is then incorporated at each denoising step, guiding the model in aligning the output image with the specific details outlined in the prompt.
@@
In particular, denoising diffusion implicit models (DDIM) sampling song2020denoising uses the predicted noise $\epsilon_{\theta}(x_{t},y,t)$ and a noise schedule represented by an array of scalars $\{\alpha_{t}\}_{t=1}^{T}$ to first predict a clean image $\hat{x}_{0}$, then makes a small step in the direction of $\hat{x}_{0}$ to obtain $x_{t-1}$. The reverse diffusion process for DDIM sampling can be formalized as follows:
$x_{t-1}=\sqrt{\alpha_{t-1}}\hat{x}_{0}+\sqrt{1-\alpha_{t-1}}\epsilon_{\theta}(x_{t},y,t)$ (2)
The training of a text-conditioned diffusion model involves minimizing a loss function that guides the model to predict the noise added to an image, given both the noisy image $x_{t}$ and the text conditioning $y$. The objective function is typically a mean squared error (MSE) between the true noise $\epsilon$ and the predicted noise $\epsilon_{\theta}(x_{t},y,t)$. The denoising model is therefore trained over the following optimization problem:
$\theta^{*}=\arg\min_{\theta}\mathbb{E}_{x,\epsilon\sim\mathcal{N}(0,I),t\sim[T]}\left[\|\epsilon-\epsilon_{\theta}(x_{t},y,t)\|^{2}\right]$ (3)
2.1.1 Style Guidance
Bansal et al. bansal2023universal propose a universal guidance algorithm that allows the reverse diffusion process to be guided by arbitrary guidance functions, including object detection, image segmentation and facial recognition. In this section, we formally describe CLIP style guidance, one of the many processes explored in the universal guidance study. We also only describe the forward guidance process. For a generalized and complete description of universal guidance process, including the backward guidance process and per-step self-recurrence, we defer the reader to the original paper.
Let $x_{c}$ denote a target image whose style we wish to replicate and $\textsf{CLIP}(\cdot)$ denote a CLIP image encoder. To incorporate style guidance, given a noisy image $x_{t}$, we first compute $\hat{x}_{0}$ using the predicted noise $\epsilon_{\theta}(x_{t},y,t)$ and Equation (1). Using $\hat{x}_{0}$, we then compute the following:
$\hat{\epsilon}_{\theta}(x_{t},y,t)=\epsilon_{\theta}(x_{t},y,t)+w\sqrt{1-\alpha_{t}}\nabla_{x_{t}}\ell_{\cos}(\textsf{CLIP}(x_{c}),\textsf{CLIP}(\hat{x}_{0}))$ (4)
where $w$ is a guidance weight parameter and $\ell_{\cos}(\cdot,\cdot)$ is the negative cosine loss.
We then perform the reverse diffusion step using the same procedure given in Equations (1) and (2), but replacing $\epsilon_{\theta}(x_{t},y,t)$ with $\hat{\epsilon}_{\theta}(x_{t},y,t)$. Using this reverse diffusion procedure, we generate a high-quality image $x_{0}$ that aligns with the text conditioning $y$ while also optimizing the cosine similarity between the CLIP embeddings of $x_{c}$ and $x_{0}$, resulting in $x_{0}$ having the stylistic characteristics of $x_{c}$.
2.1.2 Textual Inversion
Textual Inversion (TI) gal2023an is an adaptation technique that enables personalization using a small dataset of typically only 3-5 images. This approach essentially learns a new token that encapsulates the semantic meaning of the training images, allowing the model to associate specific visual features with a custom token.
To achieve this, TI trains a new token embedding, denoted as $u$, which represents a placeholder token, often denoted as $S$. During training, images are conditioned of phrases such as “A photo of $S$” or “A painting in the style of $S$”. However, unlike the fixed embeddings of typical tokens $v_{i}$, $u$ is a learnable parameter. Let $y_{u}$ denote the text conditioning vector resulting from a prompt containing the token $S$. Through gradient descent, TI minimizes the diffusion model loss with respect to $u$, while keeping the diffusion model parameters $\theta$ fixed, iteratively refining this embedding to capture the unique characteristics of the training images. The resulting optimal embedding $u_{*}$ is formalized as follows:
@@ 2.2 Differential Privacy
In this work, we adopt differential privacy (DP) dwork2006calibrating ; dwork2006differential as our privacy framework. Over the past decade, DP has become the gold standard for privacy protection in both research and industry. It measures the stability of a randomized algorithm concerning changes in an input instance, thereby quantifying the extent to which an adversary can infer the existence of a specific input based on the algorithm’s output.
Definition 2.1 ((Approximate) Differential Privacy)
@@
To achieve DP, the Gaussian mechanism is often applied dwork2014algorithmic ; balle2018improving , adding Gaussian noise scaled by the sensitivity of the function $f$ and privacy parameters $\varepsilon$ and $\delta$. Specifically, noise with standard deviation $\sigma=\frac{\Delta_{f}\sqrt{2\ln(1.25/\delta)}}{\varepsilon}$ is added to the output [^1] In practice, we use numerical privacy accountant such as balle2018improving ; mironov2017renyi to calibrate the noise. balle2018improving , where $\Delta_{f}$ represents $\ell_{2}$-sensitivity of the target function $f(\cdot)$. When the context is clear, we may omit the subscript $f$. This mechanism enables a smooth privacy-utility tradeoff and is widely used in privacy-preserving machine learning, including in DP-SGD abadi2016deep , which applies Gaussian noise during model updates to achieve DP.
2.2.1 Privacy Amplification by Subsampling
Subsampling is a standard technique in DP, where a full dataset of size $n$ is first subsampled to $m$ records without replacement (typically with $m\ll n$), and the privatization mechanism (such as the Gaussian mechanism) is then applied. Specifically, if a mechanism provides $(\varepsilon,\delta)$-DP on a dataset of size $m$, it achieves $(\varepsilon^{\prime},\delta^{\prime})$-DP on the subsampled dataset, where $\delta^{\prime}=\frac{m}{n}\delta$ and
@@ 2.2.1 Privacy Amplification by Subsampling
This result is well-known (see, for example, (steinke2022composition, , Theorem 29)). Notably, if the DP mechanism is Gaussian, even tighter amplification bounds can be applied mironov2017renyi .

### v2

@@ Abstract
We introduce a novel method for adapting diffusion models under differential privacy (DP) constraints, enabling privacy-preserving style and content transfer without fine-tuning model weights. Traditional approaches to private adaptation, such as DP-SGD, incur significant computational and memory overhead when applied to large, complex models. In addition, when adapting to small-scale specialized datasets, DP-SGD incurs large amount of noise that significantly degrades the performance. Our approach instead leverages an embedding-based technique derived from Textual Inversion (TI) and adapted with differentially private mechanisms. We apply TI to Stable Diffusion for style adaptation using two private datasets: a collection of artworks by a single artist and pictograms from the Paris 2024 Olympics. Experimental results show that the TI-based adaptation achieves superior fidelity in style transfer, even under strong privacy guarantees.
Machine Learning, ICML
Figure 1: Overview of DPAgg-TI. We first apply Textual Inversion to extract embeddings for each image in the private dataset. These embeddings are then aggregated with differentially private mechansim, incorporating subsampling to produce a private embedding $u^{\star}$. Finally, images are generated using the corresponding token <$S^{\star}$>.
@@ 1 Introduction
In recent years, diffusion models (Ho et al., 2020; Song et al., 2021b), particularly latent diffusion models (Rombach et al., 2022), have spearheaded high quality text-to-image generation, and have been widely adopted by researchers and the general public alike. Trained on massive datasets like LAION-5B (Schuhmann et al., 2022), these models have developed a broad understanding of visual concepts, enabling new creative and practical applications. Notably, tools such as Stable Diffusion (Rombach et al., 2022; Podell et al., 2023) have been made readily accessible for general use. Building on this foundation, efficient adaptation methods such as parameter efficient fine-tuning (PEFT) methods (Hu et al., 2022; von Platen et al., 2022; Ruiz et al., 2023), guidance based approaches (Ho & Salimans, 2021; Kim et al., 2022; Bansal et al., 2024), and pseudo-word generation (Gal et al., 2023) enable users to leverage this extensive pretraining for customizing models that can specialize on downstream tasks with smaller datasets.
@@
The rapid adoption of diffusion models has raised significant privacy and legal concerns. These models are vulnerable to privacy attacks, such as membership inference (Duan et al., 2023), where attackers determine if a specific data point was used for training, and data extraction (Carlini et al., 2023), which enables reconstruction of training data. This risk is amplified during fine-tuning on smaller, domain-specific datasets, where each record has a greater impact. Additionally, reliance on large datasets scraped without consent raises copyright concerns (Vyas et al., 2023), as diffusion models can reproduce original artworks without credit or compensation. These issues highlight the urgent need for privacy-preserving technologies and clearer ethical and legal guidelines for generative models.
@@
Differential privacy (DP) (Dwork et al., 2006; Dwork, 2006) is a widely adopted approach to addressing these challenges, where controlled noise is added during training or inference to prevent information leakage from individual data points while still enabling the model to learn effectively from the overall dataset. One standard approach for ensuring DP in deep learning is Differentially Private Stochastic Gradient Descent (DP-SGD) (Abadi et al., 2016), which modifies traditional SGD by adding noise to clipped gradients.
@@
However, applying DP-SGD to train diffusion models poses several challenges. It introduces significant computational and memory overhead due to per-sample gradient clipping (Hoory et al., 2021), which is essential for bounding gradient sensitivity (Dwork et al., 2006; Abadi et al., 2016). DP-SGD is also incompatible with batch-wise operations like batch normalization, as these link samples and hinder sensitivity analysis. Furthermore, training large models with DP-SGD often leads to substantial performance degradation, particularly under realistic privacy budgets since the required noise scales with the gradient norm. Consequently, existing diffusion models trained with DP-SGD are limited to relatively small-scale images. (Dockhorn et al., 2023; Ghalebikesabi et al., 2023).
@@
As a result, recent research has focused on privacy-preserving strategies for adapting diffusion models without the need for full DP-SGD training. One approach adapts large, publicly pre-trained models to new domains under DP constraints, leveraging their representational strengths while reducing computational and memory costs (Ghalebikesabi et al., 2023). Similarly, PEFT methods like DP-LoRA (Yu et al., 2022) fine-tune a small subset of parameters, enabling efficient adaptation with lower privacy costs. Methods like DP-RDM (Lebensold et al., 2024) avoid direct model updates by using retrieval mechanisms that condition image generation on private data retrieved during inference. However, these alternate approaches often fail to capture the detail of style, underscoring the challenges of balancing privacy, efficiency, and generative performance.
Independent of privacy concerns, Textual Inversion (TI) (Gal et al., 2023) provides an effective method for adapting diffusion models to specific styles or content without modifying the model. Instead, TI learns an external embedding vector that captures the style or content of a target image set, which is then incorporated into text prompts to guide the model’s outputs. A key advantage of TI is its ability to compress a style into a compact vector, reducing computational and memory demands while simplifying the application of privacy-preserving mechanisms, as privacy constraints can be applied directly to embeddings rather than the full model. Additionally, since TI avoids direct model optimization, it remains efficient and compatible with DP constraints on smaller datasets.
In this work, we propose a novel privacy-preserving adaptation method for smaller datasets, leveraging TI to avoid the extensive model updates required by DP-SGD or DP-PEFT approaches. Standard TI inherently compresses the target dataset into a single, low-dimensional vector, providing some obfuscation benefits, but it does not offer formal privacy guarantees. To address this limitation, we introduce a private variant of TI, called Differentially Private Aggregation via Textual Inversion (DPAgg-TI) and summarize it in Figure 1. Our method decouples interactions among samples by learning a separate embedding for each target image, which are then aggregated into a noisy centroid. This approach ensures efficient and secure adaptation to private datasets.
Our experiments demonstrate the effectiveness of DPAgg-TI, showing that TI remains robust in preserving stylistic fidelity even under privacy constraints. Applying our method to a private artwork collection by @eveismyname and Paris 2024 Olympics pictograms (Paris 2024, ), we show that DPAgg-TI captures nuanced stylistic elements while ensuring privacy. We observe a trade-off between privacy (controlled by DP parameter $\varepsilon$) and image quality: lower $\varepsilon$ reduces fidelity but maintains the target style under moderate noise. Subsampling further amplifies privacy by reducing sensitivity to individual data points, mitigating noise impact on image quality. This framework enables privacy-preserving adaptation of diffusion models to new styles and domains while protecting sensitive data.
Our contributions can be summarized as follows:
  • We propose DPAgg-TI that ensures privacy by learning separate embeddings for individual images and aggregating them into a noisy centroid.
  • Our approach enables style adaptation without extensive model updates, reducing computational overhead while preserving privacy.
  • We analyze the trade-off between privacy and image quality, showing that moderate noise maintains stylistic fidelity while protecting sensitive data.
  • We validate our method on diverse datasets, demonstrating its effectiveness in capturing stylistic elements under privacy constraints.
2 Preliminaries and Related Work
@@ 2.1 Diffusion Models
Diffusion models (Ho et al., 2020; Song et al., 2021b, a; Rombach et al., 2022) leverage an iterative denoising process to generate high-quality images that align with a given conditional input from random noise. In text-to-image generation, this conditional input is based on a textual description (a prompt) that guides the model in shaping the image to reflect the content and style specified by the text. To convert the text prompt into a suitable conditional format, it is first broken down into discrete tokens, each representing a word or sub-word unit. These tokens are then converted into a sequence of embedding vectors $v_{i}$ that encapsulate the meaning of each token within the model’s semantic space. Next, these embeddings pass through a transformer text encoder, such as CLIP (Radford et al., 2021), outputting a single text-conditional vector $y$ that serves as the conditioning input. This vector $y$ is then incorporated at each denoising step, guiding the model to align the output image with the specific details outlined in the prompt.
@@
The objective function for a text-conditioned diffusion model, given both the noisy image $x_{t}$ and the text conditioning $y$, is typically a mean squared error (MSE) between the true noise $\epsilon$ and the predicted noise $\epsilon_{\theta}(x_{t},y,t)$. The denoising model is therefore trained over the following optimization problem:
$\theta^{*}=\arg\min_{\theta}\mathbb{E}_{x,\epsilon\sim\mathcal{N}(0,I),t\sim[T]}[\|\epsilon-\epsilon_{\theta}(x_{t},y,t)\|^{2}].$ (1)
Textual Inversion.
Textual Inversion (TI) (Gal et al., 2023) is an adaptation technique that enables personalization using a small dataset of typically 3-5 images. This approach essentially learns a new token that encapsulates the semantic meaning of the training images, allowing the model to associate specific visual features with a custom token.
To achieve this, TI trains a new token embedding, denoted as $u$, representing a placeholder token, denoted as $S$. During training, images are conditioned on phrases such as “A photo of $S$” or “A painting in the style of $S$”. However, unlike the fixed embeddings of typical tokens $v_{i}$, $u$ is a learnable parameter. Let $y_{u}$ denote the text conditioning vector resulting from a prompt containing the token $S$. Through gradient descent, TI minimizes the diffusion model loss given in (1) with respect to $u$, while keeping the diffusion model parameters $\theta$ fixed, iteratively refining this embedding to capture the unique characteristics of the training images. The resulting optimal embedding $u_{*}$ is formalized as follows:
@@ 2.2 Differential Privacy
In this work, we adopt differential privacy (DP) (Dwork et al., 2006; Dwork, 2006) as our privacy framework. Over the past decade, DP has become the gold standard for privacy protection in both research and industry. It measures the stability of a randomized algorithm with respect to changes in an input instance, thereby quantifying the extent to which an adversary can infer the existence of a specific input based on the algorithm’s output.
Definition 2.1 ((Approximate) Differential Privacy).
@@
To achieve DP, the Gaussian mechanism is often applied (Dwork et al., 2014; Balle & Wang, 2018), adding Gaussian noise scaled by the sensitivity of the function $f$ and privacy parameters $\varepsilon$ and $\delta$. Specifically, noise with standard deviation $\sigma=\frac{\Delta_{f}\sqrt{2\ln(1.25/\delta)}}{\varepsilon}$ is added to the output [^1] In practice, we use numerical privacy accountant such as (Balle & Wang, 2018; Mironov, 2017) to calibrate the noise. (Balle & Wang, 2018), where $\Delta_{f}$ represents $\ell_{2}$-sensitivity of the target function $f(\cdot)$. When the context is clear, we may omit the subscript $f$. This mechanism enables a smooth privacy-utility tradeoff and is widely used in privacy-preserving machine learning, including in DP-SGD (Abadi et al., 2016), which applies Gaussian noise during model updates to achieve DP.
Privacy Amplification by Subsampling.
Subsampling is a standard technique in DP, where a full dataset of size $n$ is first subsampled to $m$ records without replacement (typically with $m\ll n$) before the privatization mechanism (such as the Gaussian mechanism) is applied. Specifically, if a mechanism provides $(\varepsilon,\delta)$-DP on a dataset of size $m$, it achieves $(\varepsilon^{\prime},\delta^{\prime})$-DP on the subsampled dataset, where $\delta^{\prime}=\frac{m}{n}\delta$ and
@@
This result is well-known (Steinke (2022, Theorem 29)), with tighter amplification bounds available for Gaussian mechanisms (Mironov, 2017).